# On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation

### v1

@@ 4 Case Study
This section presents a case study of the proposed bi-objective optimization problem ($\mathbf{P}_{3}$). The case study demonstrates the applicability of the proposed framework using a realistically sized constellation and representative satellite specifications. We first establish and analyze a baseline scenario, then extend the analysis through a parametric study by varying the OOS-related parameters.
@@ 4.1 Baseline Scenario
We established the baseline scenario using the constellation’s system parameters, launch service parameters, and OOS parameters, which are summarized in Tables 5, 6, and 7, respectively. The reference annual maintenance cost, $\mathrm{AMC}_{\mathrm{ref}}$, was assessed using the strategy proposed by Jakob et al. [4], which has the same supply chain as the one proposed here but without incorporating OOS, and uses the same constellation’s system and launch service parameters. The resulting optimal solution is presented in Table 8. The optimal maintenance cost is \$925.1M satisfying the service level over requirements. In addition to the scenario parameters, the decision variables and their range are summarized in Table 9.
@@
Table 9: Decision variables and their ranges for baseline scenario
Decision Variable, Unit Notation Range
$\mathbf{x}_{\mathrm{con}}$
Reorder point of in-plane spares, satellites $s$ $\left[1,20\right]$
Reorder point of parking spares, - $k_{s}$ $\left[1,20\right]$
Order quantity of parking spares, - $k_{Q}$ $\left[1,Q_{\max}\right]$
Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
$\mathbf{x}_{\mathrm{oos}}$
@@
The obtained optimal solutions are scattered in the objective space in Fig. 4. The $x$-axis and $y$-axis in Fig. 4 represent the first objective (i.e., the constellation operator’s annual maintenance cost, $\mathrm{AMC}$) and the second objective (i.e., the OOS provider’s annual profit $\mathrm{AP}$), respectively. Since the first objective is minimized and the second is maximized, the utopia point is located at the upper-left corner of the figure, marked by a red star.
The values and ranges of the decision variables, along with the resulting parameters of the optimal solutions, are summarized in Table 10. Interestingly, all solutions share identical or very similar values for the individual components of $\mathbf{x}_{\mathrm{con}}$, decision variables for the constellation operator. Notably, the optimal value of $N_{\mathrm{oos}}$ is 4, its upper bound, suggesting that once OOS is adopted, performing as many OOS operations as possible is preferable to alternative maintenance policies in this bi-objective optimization framework with the given parameters. Consequently, $T_{d}$ approaches the value of $T_{\mathrm{life}}$, 30 years.
In contrast, the OOS provider’s decision variables vary significantly, especially $p_{\mathrm{oos}}$, determining the position of each solution in the objective space. Specifically, the decisions are significantly influenced by the OOS price: as the price increases, the constellation operator’s expenditure rises, while the OOS provider’s profitability improves. On the other hand, the MTTR value is at its worst-case (upper-bound) value of 12, minimizing the OOS provider’s cost associated with service operations.
Compared with the optimal without-OOS strategy presented in Table 8, the launch cost is reduced from \$536M to \$402.4M (approximately 25% reduction), and the manufacturing cost decreases from \$160M to \$120.1M (also about 25% reduction). These reductions result from leveraging OOS, represented by $\gamma_{0}=75\%$, indicating that only 75% of failures are handled by newly launched satellites. These values remain consistent across all Pareto-optimal solutions since the constellation operator’s maintenance strategy remains unchanged, and the responsiveness of the OOS service varies only within a narrow range. Thus, the annual OOS cost, ranging from \$48.1M to \$182.4M, determines the cost-effectiveness of the OOS-integrated strategy. The Pareto-optimal solutions form a line with a slope of 1, as both objectives are linear with respect to $p_{\mathrm{oos}}$ and share the same proportional coefficient, $n_{\mathrm{oos}}N_{\mathrm{oos}}$—the total annual number of OOS service calls.
Overall, once the OOS technologies, service types, and target spacecraft subsystems are defined, the balance between the constellation operator’s expenditure and the OOS provider’s profitability across the non-dominated solutions depends primarily on the service charge. In contrast, the constellation operator’s optimal maintenance strategy remains unchanged under varying prices.
@@ 4.2 Parametric Study
[…]
@@ 4.2.1 Serviceable Failures Ratio: $r_{\mathrm{oos}}$
We first investigate the effect of the serviceable-failure ratio ($r_{\mathrm{oos}}$). In Instances 1-1 and 1-2, the $r_{\mathrm{oos}}$ values are set as 0.5 and 0.1, respectively. The results are shown in Fig. 5(a) and Table 12.
@@
From Fig. 5(a), we observed that increasing and decreasing $r_{\mathrm{oos}}$ shift the Pareto-optimal solutions toward the upper-left and lower-right directions, respectively—corresponding to movements toward and away from the utopia point. Moreover, the Pareto fronts of all instances appear nearly parallel. This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance, although differences among instances exist, as shown in Table 12.
Instance 1-1 has $N_{\mathrm{oos}}$ fixed at its upper bound, whereas Instance 1-2 has $N_{\mathrm{oos}}$ values of 3 and 4, slightly detached from the upper bound. However, despite the variation in $N_{\mathrm{oos}}$, the $\gamma_{0}$ value of Instance 1-2 remains at 0.9. This is because $r_{\mathrm{oos}}=0.1$ is sufficiently small that changes in $N_{\mathrm{oos}}$ between 3 and 4 have only a minimal effect on $\gamma_{0}$ (Eq. (9)).
Interestingly, although the OOS provider’s allowable decision ranges—those of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$—remain almost the same as in the baseline scenario, the ranges of the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, differ significantly. This implies that the constellation operator must reconsider its strategy carefully when $r_{\mathrm{oos}}$ changes substantially to maintain non-dominated solutions, even though the allowable responsiveness and OOS charge remain nearly unchanged.
By examining the cost components and $\gamma_{0}$, we found that the deployment of new satellites decreases as $r_{\mathrm{oos}}$ decreases. Specifically, from Instance 1-1 to Instances 0 and 1-2, the values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$ increase from \$283.7M to \$482.4M, from \$82.6M to \$144.0M, and from 51.6% to 90.0%, respectively, a natural outcome of the reduced contribution of OOS services.
@@
Next, we conducted experiments by varying the shift parameters of the cost-responsiveness function in Eq. (54), $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$. The resulting solutions are illustrated in Figs. 5(b) & 5(c), and the detailed values and ranges of the decision variables and result parameters are provided in Table 13.
@@
We first observed from Fig. 5(b) that increasing and decreasing $c_{\mathrm{oos},\min}$ shift the Pareto-optimal solutions toward the lower-right and upper-left directions, respectively. Figure 5(c) shows that variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have the opposite effect, but its impact is insignificant. This is because the range of optimal $\mu_{\mathrm{oos}}$ lies in the lower-right tail of the cost-responsiveness curve shown in Fig. 3, where translation along the $x$-axis has only a limited effect on $c_{\mathrm{oos}}$.
Similar to the results of the previous instances, the Pareto fronts of all cases appear nearly parallel. This occurs because the solutions share the same (or very similar) constellation operator maintenance strategies and comparable levels of responsiveness within each instance. In addition, we can see that the constellation operator’s decision variables remain the same as those in the baseline scenario. However, the minimum values of the optimal $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ shift in Instance 2-1 and Instance 2-2, respectively, due to the translation of the cost-responsiveness curve along the $y$- and $x$-axes.
Because the constellation operator’s decision variables remain identical to those in the baseline scenario, the dependence on OOS does not change, as reflected by the unchanged values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$. Nevertheless, variations in service charge and responsiveness lead to changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$, through slight adjustments in $C_{\mathrm{hold}}$ and more substantial changes in $C_{\mathrm{oos}}$.
@@ 4.2.3 Shape Parameters of Cost-Responsiveness Function: $\alpha_{1}$ & $\alpha_{2}$
Lastly, we vary the shape parameters of the cost-responsiveness function, $\alpha_{1}$ and $\alpha_{2}$. The resulting solutions are presented in Figs. 5(d) & 5(e) and Table 14.
@@
From Figs. 5(d) & 5(e), we observed that an increase in $\alpha_{1}$ shifts the Pareto-optimal solutions toward the lower-right direction, whereas a change in $\alpha_{2}$ produces the opposite shift. The trend associated with $\alpha_{1}$ is intuitive, as increasing $\alpha_{1}$ raises the OOS cost for a given $\mu_{\mathrm{oos}}$. However, the effect of $\alpha_{2}$ depends on the value of $\mu_{\mathrm{oos}}$: when $\mu_{\mathrm{oos}}$ lies near the lower-right tail of the cost-responsiveness curve (i.e., for small $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ increases the OOS cost; conversely, when $\mu_{\mathrm{oos}}$ lies near the upper-left tail (i.e., for large $\mu_{\mathrm{oos}}$), increasing $\alpha_{2}$ decreases the OOS cost. Because the current Pareto-optimal solutions occupy the region near the lower-right tail of the cost-responsiveness curve, variations in $\alpha_{2}$ move the overall solutions in the opposite direction of $\alpha_{1}$.
As with the previous instances, the Pareto fronts across all cases remain nearly parallel, for the same underlying reasons discussed earlier. Notably, the minimum values of $p_{\mathrm{oos}}$ vary between instances, whereas the maximum values remain almost constant, consistent with the previously observed patterns. In contrast to earlier cases, the maximum values of $\mu_{\mathrm{oos}}^{-1}$ change when either $\alpha_{1}$ decreases or $\alpha_{2}$ increases, as these adjustments make the cost-responsiveness function more favorable in shape, allowing greater responsiveness to be achieved with only a marginal increase in OOS cost.
Consistent with Instances 2-X and 3-X, the constellation operator’s decision variables are identical to those in the baseline scenario. As a result, the level of dependence on OOS remains unchanged, as reflected in the constant values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$. However, variations in the service charge and responsiveness still lead to corresponding changes in the two objectives, $\mathrm{AMC}$ and $\mathrm{AP}$.
4.3 Results Summary and Strategic Implications
Through the case study, we observed Pareto-efficient solutions of $\mathbf{P}_{3}$ with variations in OOS-related parameters ($r_{\mathrm{oos}}$, $c_{\mathrm{oos},\min}$, $\mu_{\mathrm{oos},\mathrm{ideal}}$, $\alpha_{1}$, $\alpha_{2}$), while keeping the constellation and launch service parameters fixed. The summary of the results and their implications is provided below.
Given the constellation and launch service parameters, as well as the fraction of serviceable failures ($r_{\mathrm{oos}}$), Pareto-optimal solutions share a common maintenance strategy adopted by the constellation operator. This observation suggests that, to remain Pareto-efficient, there exists a dominant maintenance strategy for the satellite constellation. However, each Pareto-efficient solution differs in the OOS provider’s decision variables, particularly the OOS price $p_{\mathrm{oos}}$, which governs the location of the Pareto-optimal solution in the objective space. As $p_{\mathrm{oos}}$ increases, the constellation operator’s annual maintenance cost ($\mathrm{AMC}$) rises, while the OOS provider’s annual profit ($\mathrm{AP}$) also increases. Therefore, $p_{\mathrm{oos}}$ acts as a balancing parameter between the constellation operator’s expenditure and the OOS provider’s profits. Although the responsiveness of OOS ($\mu_{\mathrm{oos}}$) varies among different Pareto-optimal solutions, the variations remain within a limited range and have only minor effects on the results, since they lie within the right-bottom tail of the cost-responsiveness curve shown in Fig. 3. Consequently, the Pareto-optimal solutions form a linear trend with a slope of 1, cut by lines of $\mathrm{AP}=0$ (as the OOS cost cannot exceed the OOS price) and $\mathrm{AMC}=\mathrm{AMC}_{\mathrm{ref}}$ (due to the maintenance cost threshold set by the constellation operator). The slope of 1 is attributed to variations in $p_{\mathrm{oos}}$ affecting $\mathrm{AMC}$ linearly via Eq. (34) and $\mathrm{AP}$ linearly via Eq. (53), both with coefficients of $n_{\mathrm{oos}}N_{\mathrm{plane}}$.
Among the OOS-related parameters, the only factor that alters the constellation operator’s maintenance strategy is $r_{\mathrm{oos}}$, the serviceable failure ratio. When $r_{\mathrm{oos}}$ is sufficiently high, Pareto-optimal solutions consistently yield the maximum value of $N_{\mathrm{oos}}$, indicating that the constellation operator fully utilizes the OOS capability. However, as $r_{\mathrm{oos}}$ decreases significantly, $N_{\mathrm{oos}}$ may deviate from its upper bound, since it has little impact on the fraction of failures replaced by newly launched satellites, represented by the ratio $\gamma_{0}$ (or equivalently, the fraction of failed satellites serviced by OOS, $1-\gamma_{0}$). As a result, increasing $r_{\mathrm{oos}}$ raises the fraction of failures handled by OOS, thereby increasing the proportion of the total annual maintenance cost attributed to OOS. Conversely, when $r_{\mathrm{oos}}$ decreases, the opposite trend occurs.
Regarding the cost-responsiveness function, variations in the shift parameters—the ideal OOS cost ($c_{\mathrm{oos},\min}$) and responsiveness ($\mu_{\mathrm{oos},\mathrm{ideal}}$)—minimally affect the decision variables of the Pareto-optimal solutions. Since the Pareto-optimal solutions lie within the right-bottom tail of the cost-responsiveness curve, changes in $c_{\mathrm{oos},\min}$ directly shift the OOS cost, whereas variations in $\mu_{\mathrm{oos},\mathrm{ideal}}$ have negligible influence on the overall cost and results. In contrast, variations in the shape parameters $\alpha_{1}$ and $\alpha_{2}$ significantly affect the efficient ranges of $p_{\mathrm{oos}}$ and $\mu_{\mathrm{oos}}$ by reshaping the overall curvature of the cost-responsiveness curve. Increasing $\alpha_{1}$ raises the overall cost for a given responsiveness, shifting the Pareto-optimal solutions toward the degrading direction. Conversely, because the Pareto-optimal solutions are located near the right-bottom tail of the curve, increasing $\alpha_{2}$ moves the Pareto-optimal solutions toward the improving direction.

### v2

@@ 5 Case Study
This section presents a case study of the proposed bi-objective optimization problem ($\mathbf{P}$). The case study demonstrates the applicability of the proposed framework using a realistically sized constellation and representative satellite specifications. We first establish and analyze a baseline scenario, then extend the analysis through a parametric study by varying the OOS-related parameters.
@@ 5.1 Baseline Scenario
We established the baseline scenario using the constellation’s system parameters, launch service parameters, and OOS parameters, which are summarized in Tables 7, 8, and 9, respectively. The decision variables and their corresponding ranges are listed in Table 10. The constellation and launch service parameters in Tables 7 and 8 are adopted from prior studies [31, 32] to reflect realistic, real-world-scale systems. By contrast, the OOS-related parameters cannot yet be assigned fully realistic values, because the relevant service capabilities and market conditions remain uncertain. Instead, they are specified within plausible ranges based on the existing literature, and their effects on the results are further examined in the parametric study by varying their values. In particular, the minimum MTTR is set to $2$ weeks (i.e., $\mu_{\mathrm{oos}}=1/2$), which lies within a reasonable range according to prior studies on OOS responsiveness [50, 77, 51]. The serviceable failure ratio, $r_{\mathrm{oos}}$, is set to 2525%, considering the maximum fraction of failures that could be induced by a single spacecraft subsystem failure, namely the thruster subsystem, as discussed by Saleh and Castet [84]. The minimum OOS cost is set equal to the satellite production cost, $c_{\mathrm{sat}}$. The shape parameters, $\alpha_{1}$ and $\alpha_{2}$, are both set to 1.
@@
Table 10: Decision variables and their ranges for baseline scenario
Decision Variable, Unit Notation Range
Reorder point of in-plane spares, satellites $s$ $\left[1,20\right]$
Reorder point of parking spares, - $k_{s}$ $\left[1,20\right]$
Order quantity of parking spares, - $k_{Q}$ $\left[1,Q_{\max}\right]$
Maximum number of services, - $N_{\mathrm{oos}}$ $\left[1,4\right]$
[…]
@@
Using these parameter settings, the bi-objective optimization problem ($\mathbf{P}$) was solved using NSGA-II [83], implemented in Python with the pymoo package [85]. The algorithm was run with a population size of 400 for 200 generations, while all other algorithm parameters were kept at their default values. To assess the reliability of the results, we performed three independent runs with different random seeds. Because the ranges of the obtained solutions were sufficiently similar across the three runs in both the decision and objective spaces, we retained the solution set generated using the first random seed. The obtained optimal solutions are scattered in the objective space in Fig. 5. The $x$-axis and $y$-axis in Fig. 5 represent the first objective (i.e., the constellation operator’s annual maintenance cost, $\mathrm{AMC}$) and the second objective (i.e., the OOS provider’s annual profit $\mathrm{AP}$), respectively. Since the first objective is minimized and the second is maximized, the utopia point is located at the upper-left corner of the figure, marked by a red star.
@@
The values and ranges of the decision variables, together with the resulting performance measures of the Pareto-optimal solutions, are summarized in Table 12. Figure 5 shows that, under the baseline setting, the Pareto-optimal solutions form an almost linear frontier spanning $\mathrm{AMC}=790.6$$925.0$ \$/year and $\mathrm{AP}=0$$134.5$ \$/year. This frontier is bounded by the OOS provider’s non-negative profit condition (Eq. (51)) and the constellation operator’s annual maintenance cost threshold, $\mathrm{AMC}_{\mathrm{ref}}$ (Eq. (52)). In addition, the constraints in Eqs. (41)–(43) and Eq. (50) are effectively binding along the Pareto frontier. Although they are not satisfied at equality for every solution, the remaining slack is small and is attributable to the integrality of the decision variables. Relative to the benchmark solution without OOS in Table 11, the constellation operator can reduce annual maintenance cost by as much as \$134.5M/year (14.5%) while maintaining the required service levels. Moreover, the Pareto-optimal set is generated by an almost invariant maintenance strategy on the constellation operator side: $s=3$, $Q=4$, $k_{s}=6$, $k_{Q}=10$, $N_{\mathrm{parking}}=7$, $h_{\mathrm{parking}}\in[674.3,700.4]$ km, and $N_{\mathrm{oos}}=4$. This indicates that, in the baseline environment, the location of a Pareto solution is governed primarily by the OOS provider’s decision rather than by multiple qualitatively different maintenance strategies for the constellation operator.
Compared with the optimal strategy without OOS, the OOS-integrated solutions reduce annual launch cost from \$536.0M/year to \$402.4M/year and annual manufacturing cost from \$160.0M/year to \$120.1M/year, both reductions of approximately 25%. These savings are partially offset by annual OOS expenditures of \$48.1M/year to \$182.4M/year. Since $\gamma_{0}=75.1\%$, approximately 24.9% of failures are handled through OOS rather than by newly launched replacements. Along the Pareto frontier, the main variation comes from the OOS price $p_{\mathrm{oos}}$, whereas the optimal OOS MTTR remains near its upper bound, $\mu_{\mathrm{oos}}^{-1}=10.6$$12.0$ weeks. This implies that, under the assumed cost-responsiveness relationship, the provider maximizes profit by selecting relatively slow but inexpensive service. Accordingly, the Pareto frontier is nearly linear, indicating that movement along it primarily redistributes value between the constellation operator and the OOS provider, rather than substantially altering the physical structure of the maintenance system.
@@ 5.2 Parametric Study
We next examine how the Pareto-optimal solutions change when the OOS-related parameters are perturbed around the baseline scenario. Specifically, we vary the serviceable-failure ratio $r_{\mathrm{oos}}$, the lower bound of OOS cost $c_{\mathrm{oos},\min}$, the ideal responsiveness parameter $\mu_{\mathrm{oos},\mathrm{ideal}}$, and the two shape parameters $\alpha_{1}$ and $\alpha_{2}$, as summarized in Table 13. All optimization runs were performed using the same algorithm settings as in the baseline scenario. To avoid repetitive interpretation, the results are discussed in terms of three questions: how each parameter shifts the Pareto frontier, whether it changes the constellation operator’s preferred maintenance strategy, and what mechanism explains the observed change.
@@ 5.2.1 Serviceable Failures Ratio: $r_{\mathrm{oos}}$
Among the tested parameters, the serviceable-failure ratio $r_{\mathrm{oos}}$ has the strongest structural effect, because it directly determines the share of failures that can be covered by the production-based replenishment to the recovery from OOS. As shown in Fig. 6(a) and Table 14, increasing $r_{\mathrm{oos}}$ from $0.1$ to $0.5$ shifts the Pareto frontier markedly toward the utopia point: the minimum $\mathrm{AMC}$ decreases from \$873.4M/year to \$661.9M/year, while the maximum $\mathrm{AP}$ increases from \$51.7M/year to \$263.1M/year. This improvement is accompanied by substantial reductions in annual launch and manufacturing costs, from \$482.4M/year to \$283.7M/year and from \$144.0M/year to \$82.6M/year, respectively. At the same time, $\gamma_{0}$ decreases from 90.0% to 51.6%, implying that the fraction of failures handled through OOS increases from 10.0% to 48.4%.
@@
Unlike the other OOS-related parameters, $r_{\mathrm{oos}}$ also changes the constellation operator’s preferred maintenance configuration. When $r_{\mathrm{oos}}=0.5$, the optimal value of $N_{\mathrm{oos}}$ remains fixed at its upper bound. In contrast, when $r_{\mathrm{oos}}=0.1$, $N_{\mathrm{oos}}$ takes values of 3 or 4, and the optimal stocking decisions move closer to those of the without-OOS benchmark in Table 11: $s$, $Q$, $k_{Q}$, and $N_{\mathrm{parking}}$ coincide with the benchmark values, while $k_{s}$ and $h_{\mathrm{parking}}$ differ only modestly. This result indicates that $r_{\mathrm{oos}}$ is the key parameter determining whether OOS functions merely as an economic supplement or becomes a structurally important component of the maintenance architecture.
@@
We next examine the shift parameters of the cost-responsiveness function, $c_{\mathrm{oos},\min}$ and $\mu_{\mathrm{oos},\mathrm{ideal}}$. The resulting solutions are illustrated in Figs. 6(b) and 6(c), while the detailed values and ranges of the decision variables and outcome measures are reported in Table 15. Their main effect is to shift the Pareto frontier without structural change of the constellation operator’s preferred maintenance strategy. Across Instances 2-X and 3-X, the operator-side decision variables remain identical to those of the baseline scenario, and the corresponding dependence on OOS also remains unchanged, as reflected by the constant values of $C_{\mathrm{lau}}=\mathdollar 402.4$M/year, $C_{\mathrm{manufac}}=\mathdollar 120.1$M/year, and $\gamma_{0}=75.1\%$.
@@
The effect of $c_{\mathrm{oos},\min}$ is direct and substantial. Increasing $c_{\mathrm{oos},\min}$ from \$0.25M to \$1M shifts the Pareto frontier downward and to the right, increasing the minimum $\mathrm{AMC}$ from \$770.7M/year to \$830.5M/year and reducing the maximum $\mathrm{AP}$ from \$154.4M/year to \$94.6M/year. By contrast, varying $\mu_{\mathrm{oos},\mathrm{ideal}}$ has only a minor effect: changing it from $0.25$ to $1$ week-1 changes the minimum $\mathrm{AMC}$ only from \$792.7M/year to \$789.9M/year and the maximum $\mathrm{AP}$ only from \$132.4M/year to \$135.2M/year. This limited sensitivity arises because the efficient solutions remain in the low-responsiveness region of the cost-responsiveness curve, where horizontal shifts of the asymptote have only a small influence on cost.
@@ 5.2.3 Shape Parameters of Cost-Responsiveness Function: $\alpha_{1}$ & $\alpha_{2}$
Lastly, we vary the shape parameters $\alpha_{1}$ and $\alpha_{2}$, which reshape the curvature of the cost-responsiveness function rather than shift it rigidly. The resulting solutions are presented in Figs. 6(d) & 6(e) and Table 16. Their effects are moderate and, again, do not alter the constellation operator’s maintenance strategy. Across all four instances, the operator-side decision variables remain identical to those in the baseline scenario, and the values of $C_{\mathrm{lau}}$, $C_{\mathrm{manufac}}$, and $\gamma_{0}$ therefore remain unchanged.
@@
Increasing $\alpha_{1}$ makes the cost-responsiveness curve less favorable and shifts the Pareto frontier away from the utopia point: the minimum $\mathrm{AMC}$ increases from \$786.1M/year for $\alpha_{1}=0.5$ to \$798.6M/year for $\alpha_{1}=2$, while the maximum $\mathrm{AP}$ decreases from \$139.0M/year to \$126.5M/year. The effect of $\alpha_{2}$ is reversed in the present problem because the efficient solutions lie near the low-responsiveness tail of the curve. Specifically, increasing $\alpha_{2}$ from $0.5$ to $2$ improves the frontier, reducing the minimum $\mathrm{AMC}$ from \$807.9M/year to \$781.8M/year and increasing the maximum $\mathrm{AP}$ from \$117.2M/year to \$143.2M/year. Hence, the curvature of the provider’s cost technology matters, but mainly through the economic terms of OOS rather than through a redesign of the constellation maintenance system.
5.3 Discussions
5.3.1 Main Findings and Strategic Implications
Three main findings emerge from the case study. First, under fixed constellation, launch service, and OOS-related parameters, the Pareto-optimal set within a given scenario (with fixed OOS-related parameters) is generated by an almost invariant maintenance strategy on the constellation operator’s side. In the baseline scenario, the optimal operator-side decisions remain unchanged across the Pareto frontier, while the main variation comes from the provider-side pricing decision, with only limited variation in OOS responsiveness. As a result, the baseline Pareto frontier spans approximately $(\mathrm{AMC},\mathrm{AP})=(\mathdollar 925.0\text{M/year},0)$ to $(\mathdollar 790.6\text{M/year},\mathdollar 134.5\text{M/year})$. This indicates that, relative to the benchmark solution without OOS, OOS can create up to \$134.5M/year of maintenance-cost savings for the constellation operator while still allowing non-negative profit for the OOS provider. A similar within-instance pattern is also observed across every scenario in the parametric study.
Second, OOS provides material logistical relief even before considering how value is divided between the two stakeholders. Relative to the benchmark solution without OOS, the baseline OOS-integrated solutions reduce annual launch cost from \$536.0M/year to \$402.4M/year and annual manufacturing cost from \$160.0M/year to \$120.1M/year, corresponding to reductions of approximately 25% in both components, while maintaining the required service levels of approximately 98% throughout the case study. In the baseline case, $\gamma_{0}=75.1\%$, implying that approximately 24.9% of failures are handled through OOS rather than by newly launched replacements. This shows that the value of OOS lies not only in the financial transfer between stakeholders but also in its ability to reduce dependence on new satellite production and launch. In a broader sense, this may also contribute to more sustainable use of the space environment, although such implications lie beyond the scope of the present model.
Third, among the OOS-related parameters, the serviceable-failure ratio $r_{\mathrm{oos}}$ is the only one that materially changes the constellation operator’s preferred maintenance architecture. Increasing $r_{\mathrm{oos}}$ from 0.1 to 0.5 reduces the minimum $\mathrm{AMC}$ from \$873.4M/year to \$661.9M/year and increases the maximum $\mathrm{AP}$ from \$51.7M/year to \$263.1M/year, while also shifting the operator-side decisions farther from the benchmark solution without OOS and toward stronger reliance on OOS. By contrast, variations in $c_{\mathrm{oos},\min}$, $\mu_{\mathrm{oos},\mathrm{ideal}}$, $\alpha_{1}$, and $\alpha_{2}$ mainly shift or reshape the Pareto frontier without materially changing the operator-side configuration. From a managerial standpoint, this implies that expanding the set of recoverable failure modes may be at least as important as reducing service cost or improving nominal responsiveness.
5.3.2 Limitations
These findings should be interpreted in light of several limitations. First, the cost-responsiveness relationship is imposed exogenously through a specific parametric functional form. Although its qualitative shape is broadly plausible, given the underlying effort/cost trade-off and its consistency with prior work, it may not accurately represent future OOS technologies, operational concepts, or market structures. Second, recovered satellites are assumed to return to an “as-good-as-new” state, which likely overstates the long-run benefit of repeated servicing. Third, the analysis adopts a two-stakeholder bi-objective optimization framework and therefore abstracts from bargaining, competition among OOS providers, contracting frictions, and other strategic effects that may determine which solution is realized in practice. As a result, the realized outcome need not coincide with the Pareto-optimal frontier identified by the present model and may, in practice, lie outside it.